JET COOLED ROTATIONAL SPECTRA OF LASER ABLATED URACIL AND THYMINE

M. ISABEL PENA, VANESA VAQUERO, \underline{JUAN C. LÓPEZ}, and JOSÉ L. ALONSO, Grupo de Espectroscopía Molecular (GEM), Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47005 Valladolid, Spain.

The rotational spectra of the nitrogen bases uracil^{a} (m.p. 335 ^{∘}C) and thymine^{b} (m.p. 316 ^{∘}C) have been observed in the 6-18 GHz frequency range by using a Laser-Ablation Molecular Beam Fourier Transform Microwave Spectrometer (LA-MB-FTMW).^{c} The nuclear quadrupole coupling hyperfine structure corresponding to the presence of two ^{14}N atoms has been completely resolved for both molecular systems. The rotational spectrum of all uracil ^{13}C and ^{15}N monosubstituted isotopomers in their natural abundance are being observed. Following the characterization of these isotopomers, the uracil structure in gas phase will be obtained.

a R. D. Brown, P. D. Godfrey, D. McNaughton and A. Pierlot, J. Am. Chem. Soc., 110, 2329 (1988)

b R. D. Brown, P. D. Godfrey, D. McNaughton and A. Pierlot, J. Chem. Soc. Chem. Commun., 37, (1989)

c A. Lesarri, S. Mata, J. C. López and J. L. Alonso, Rev. Sci. Instrum., 74, 4799 (2003).